# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

### v1

@@ 3.1 Dataset and Preprocessing
In the experiment, we downsampled the original high-density EEG data along the channel dimension to obtain corresponding low-resolution (LR) counterparts at different spatial resolutions, with scaling factors approximately set to 2, 4, 8, and 16. To simulate the acquisition of low-density EEG data in the real world, as Table 1 shows, we mapped each scaling factor to different numbers of channels in the EGI standard electrode systems[42], obtaining corresponding EEG electrode arrays. Thus, by using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG data that match the preset channel levels. These LR-HR data pairs at different spatial resolutions were then created for model training.
@@ 3.3.1 Comparison With Other Existing EEG SR Methods
the proposed STADMs are compared with the existing EEG spatial SR methods to demonstrate the effectiveness of the synthetic HR EEG (256 channels). We mainly quantitatively analyze the reconstruction performance differences between the proposed STADMs and CNN-based methods (Deep-CNN [17]), autoencoder-based methods [Weighted gate layer autoencoders (WGLAE) [22], Deep autoencoder (DAE) [44]], and GANs-based method (EEGSR-GAN) [23].
Fig. 2 shows quantitative comparison results between different methods. As observed, the proposed STADMs achieve the best performance, significantly outperforming other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs can maintain high signal quality and correlation. Moreover, STADMs significantly outperforms other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG spatial super-resolution methods, the proposed STADMs can effectively reconstruct high-quality high spatial resolution EEG (256 channels), achieving high spatial resolution enhancement of LR EEG.
@@
3.3.2 Comparison With Different Scale Factors
In order to evaluate the performance of the proposed STADMs in synthetic SR EEG signals from LR EEG under different scale factors, we employ four metrics for quantitative analysis of the model generative performance. Fig. 3 shows quantitative results for different scale factors. As the scale factor increases, PCC and SNR values decrease, indicating a weakening correlation between reconstructed SR EEG and ground truth; meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that STADMs’s reconstruction performance diminishes at larger scale factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scale factors, STADMs’s performance remains comparable to smaller scale factors. For example, with a scale factor of 16, the PCC value is only 3.02 % lower than at a scale factor of 8, the MAE value differs by 1.68 %, and the SNR value by 3.19 %. This indicates STADMs can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STADMs effectively enhances LR EEG spatial resolution across different scale factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
@@ 3.3.3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands.We first use the Multitaper method to extract the power spectral density (PSD) of LR EEG and corresponding synthetic SR EEG in five frequency bands, including (0.5 4 Hz), (4 8 Hz), (8 13 Hz), (13 30 Hz), and (30 40 Hz or higher) [45]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 4 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low (, , ) and high-frequency bands (, ), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.
@@ 3.4 Classification Evaluation on the Synthetic SR EEG
Enhancing the spatial resolution of EEG significantly improves the performance of EEG-based diagnostic systems. For epilepsy patients, HR EEG aids doctors in identifying abnormal waveforms, such as spikes and sharp waves during seizures. To evaluate the effectiveness of synthetic SR EEG in detecting abnormalities, we designed a binary classification experiment (normal vs. abnormal) using SR EEG. In the Localize-MI dataset, EEG signals during stimulation are labeled abnormal, and those before stimulation are labeled normal. We train the EEG-Net [46] separately on both LR EEG and the synthetic SR EEG, and conduct corresponding classification tests. This allows us to compare the performance of different spatial resolution LR EEG and synthetic SR EEG in the classification task.
Fig. 5 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STADMs outperforms LR EEG in all metrics, showing that STADMs enhances classification performance. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STADMs effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
@@
| scale factor 4 | scale factor 8 | scale factor 16 | ||
@@
We also analyze classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to n 5 features (n is the number of channels), and classification is based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scale factors. SR EEG outperforms LR EEG in all frequency bands. As the scale factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands (, , ) have worse performance than high-frequency bands (, ), but the average accuracy improvement is higher in low-frequency bands (5.06 %) compared to high-frequency bands (4.95 %). This indicates that the proposed method better preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
@@ 3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [47]. To evaluate the effectiveness of the proposed method in enhancing the spatial resolution of EEG and improving downstream task performance, we assess it using the source localization task, including both quantitative and qualitative analyses. For quantitative analysis, we use localization error [7] as the metric, quantified as the Euclidean distance between the source localization results and the true source positions, to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [48] as the source localization method, an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [49]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
Fig. 6 shows a comparison of localization error for different EEG types among all subjects in the dataset. The localization error for LR EEG is widely dispersed with higher values, indicating lower precision and greater variability. In contrast, SR EEG shows significantly lower and more concentrated localization errors, suggesting higher precision and stability. Although localization error of SR EEG is slightly higher than HR EEG, its accuracy is markedly improved, making it a viable substitute for HR EEG and reducing reliance on expensive high-density equipment. Therefore, the experimental results demonstrate that the synthetic SR EEG substantially enhances the spatial resolution and source localization accuracy of LR EEG.
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results and blue circles show the results from the LR EEG. The LR EEG results exhibit significant deviations from the ground truth, whereas SR EEG results show minimal deviation, closely matching HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
@@ 4 Discussion and Conclusion
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scale factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) At larger scale factors, the available spatial information for the model is limited, making it difficult to accurately capture the spatial relationships between channels; ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model; iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.

### v2

@@ 3.1 Dataset and Preprocessing
In the experiment, we downsampled the original HR EEG data along the channel dimension to obtain LR counterparts at different spatial resolutions, with scaling factors approximately set to 2, 4, 8, and 16. To simulate the acquisition of low-density EEG data in the real world, we mapped each scaling factor to different numbers of channels in the EGI standard electrode systems[42], as shown in Table 1. This process involved obtaining corresponding EEG electrode arrays. By using electrode systems with different channel ranks, we identified and removed the channels to be estimated, resulting in LR EEG that matched the preset channel levels. These LR-HR data pairs at different spatial resolutions were then created for model training.
@@ 3.3.1 Comparison With Other Existing EEG SR Methods
the proposed STADMs are compared with the existing EEG SR methods to demonstrate the effectiveness of the synthetic HR EEG (256 channels). We quantitatively analyze the reconstruction performance differences between the proposed STADMs and several other methods: CNN-based methods (Deep-CNN [17]), autoencoder-based methods [Weighted gate layer autoencoders (WGLAE) [22], Deep autoencoder (DAE) [44]], and GANs-based method (EEGSR-GAN) [23].
Fig. 2 shows quantitative comparison results between different methods. As observed, the proposed STADMs achieve the best performance, significantly outperforming other existing competitive methods. The highest PCC and SNR values indicate that the reconstructed results of STADMs maintain high signal quality and correlation. Moreover, STADMs significantly outperform other methods in terms of MAE and NMSE, indicating lower errors and higher accuracy in reconstructing SR EEG. Additionally, the gaps between STADMs and other methods are substantial across all quantitative metrics. Through the experimental results, it can be concluded that compared to existing EEG SR methods, the proposed STADMs effectively reconstruct high spatial resolution(256 channels) EEG, achieving significant spatial resolution enhancement of LR EEG.
@@
3.3.2 Comparison With Different scaling factors
To evaluate the performance of the proposed STADMs in generating synthetic SR EEG from LR EEG under different scaling factors, we employ four metrics for quantitative analysis of the model’s generative performance. Fig. 3 shows quantitative results for different scaling factors. As the scaling factor increases, PCC and SNR values decrease, indicating a weakening correlation between synthetic SR EEG and ground truth. Meanwhile, MAE and NMSE values increase, implying a growing difference. This suggests that the reconstruction performance of STADMs diminishes at larger scaling factors due to limited channel information, making HR EEG reconstruction challenging. However, even at larger scaling factors, the performance of STADMs remains comparable to that at smaller scaling factors. For example, with a scaling factor of 16, the PCC value is only 3.02% lower than at a scaling factor of 8, the MAE value differs by 1.68%, and the SNR value by 3.19%. This indicates that STADMs can still adequately reconstruct SR EEG signals despite limited channel information. Overall, STADMs effectively enhance LR EEG spatial resolution across different scaling factors, achieving better reconstruction at smaller scales while maintaining satisfactory performance at larger scales, demonstrating strong adaptability and robustness in handling various channel levels.
@@ 3.3.3 Qualitative results of Synthetic EEG
To observe the differences between synthetic SR EEG and LR EEG at different channel levels (16, 32, 64, and 128) more intuitively, we analyze the scalp topomaps of EEG in different frequency bands.We first use the Multitaper method to extract the power spectral density (PSD) of LR EEG and corresponding synthetic SR EEG in five frequency bands, including (0.5-4 Hz), (4-8 Hz), (8-13 Hz), (13-30 Hz), and (30-40 Hz or higher) [45]. Finally, we use the MNE tool to obtain scalp topographies in different frequency bands. Fig. 4 compares scalp topographies across different frequency bands between EEG at various channel levels and model-generated 256-channel SR EEG. Each column represents a frequency band, and each row represents a type of EEG. As spatial resolution increases, the active regions on the scalp topographies become smaller and clearer. For instance, the 16-channel LR EEG shows large, blurry active regions in both low (, , ) and high-frequency bands (, ), whereas the 256-channel SR EEG accurately localizes active sources in high-frequency bands. The synthesized SR EEG retains the frequency characteristics of the original signal across all bands, effectively reconstructing their spatial distribution. This demonstrates the strong generalization of the model across different frequency bands and its ability to stably reconstruct EEG signals. Overall, the proposed method generates SR EEG with richer spatial detail, more accurately describing and locating active regions on the scalp, whereas LR EEG results in coarse active regions, making it challenging to pinpoint active brain areas.
@@ 3.4 Classification Evaluation on the Synthetic SR EEG
Enhancing the spatial resolution of EEG significantly improves the performance of EEG-based diagnostic systems. For epilepsy patients, HR EEG aids doctors in identifying abnormal waveforms, such as spikes and sharp waves during seizures. To evaluate the effectiveness of synthetic SR EEG in detecting abnormalities, we designed a binary classification experiment (normal vs. abnormal) using SR EEG. In the Localize-MI dataset, EEG signals during stimulation are labeled abnormal, and those before stimulation are labeled normal. We trained the EEG-Net [46] separately on both LR EEG and the synthetic SR EEG and conducted corresponding classification tests. This allowed us to compare the performance of different spatial resolution LR EEG and synthetic SR EEG in the classification task.
Fig. 5 shows the comparison results of four classification metrics between LR EEG and synthetic SR EEG. The SR EEG reconstructed by STADMs outperforms LR EEG in all metrics, showing that STADMs enhance classification performance of EEG-Net. As the channel count of LR EEG increases, the classification performance improvement by SR EEG also increases, indicating better detail reconstruction. This demonstrates that STADMs effectively enhance LR EEG spatial resolution, improving downstream classification tasks. Moreover, this provides a novel approach for other EEG signal enhancement and EEG-based clinical diagnosis research.
@@
| scaling factor 4 | scaling factor 8 | scaling factor 16 | ||
@@
We also analyzed classification performance from the frequency domain perspective, calculating the PSD of LR and SR EEG in five frequency bands. Each sample corresponds to n 5 features (where n is the number of channels), and classification is based on these features. Table 2 shows classification results for LR and SR EEG across different frequency bands and scaling factors. SR EEG outperforms LR EEG in all frequency bands. As the scaling factor increases, classification accuracy decreases for both but remains higher for SR EEG. Low-frequency bands (, , ) have worse performance than high-frequency bands (, ), but the average accuracy improvement is higher in low-frequency bands (5.06%) compared to high-frequency bands (4.95%). This indicates that the proposed method better preserves low-frequency information and captures high-frequency variations, improving classifier performance in identifying abnormal EEG features.
@@ 3.5 Source Localization Evaluation on the Synthetic EEG
Source localization is a crucial task in neuroscience and clinical applications, as it reveals the active source locations within the brain [47]. To evaluate the effectiveness of the proposed method in enhancing the spatial resolution of EEG and improving downstream task performance, we assess it using the source localization task, incorporating both quantitative and qualitative analyses. For quantitative analysis, we use localization error [7] as the metric, quantified as the Euclidean distance between the source localization results and the true source positions, to compare the performance of LR EEG, SR EEG, and HR EEG. For qualitative analysis, we visualize the source localization results on brain maps to provide an intuitive display of the localization performance across different data sources. To ensure fairness, we consistently use eLORETA [48], an unconstrained linear inverse solution that infers the distribution of brain current sources from scalp-recorded EEG signals through precise regularization strategies [49]. By examining both quantitative results and visualization outcomes, we can evaluate whether the proposed method effectively enhances the spatial resolution of EEG.
Fig. 6 shows a comparison of localization errors for different EEG types among all subjects in the dataset. The localization error for LR EEG is widely dispersed with higher values, indicating lower precision and greater variability. In contrast, SR EEG shows significantly lower and more concentrated localization errors, suggesting higher precision and stability. Although the localization error of SR EEG is slightly higher than HR EEG, its accuracy is markedly improved, making it a viable substitute for HR EEG and reducing reliance on expensive high-density equipment. Therefore, the experimental results demonstrate that synthetic SR EEG substantially enhances the spatial resolution and source localization accuracy of LR EEG.
Fig. 7 shows the visualization results of different EEG data types on source localization tasks for some subjects. In these visualizations, green circles represent the stimulus source location ground-truth, red circles indicate the original HR EEG results, yellow circles denote the synthesized SR EEG results, and blue circles show the results from the LR EEG. The LR EEG results exhibit significant deviations from the ground truth, whereas SR EEG results show minimal deviation, closely matching HR EEG performance. This demonstrates that the proposed method effectively enhances spatial resolution and signal quality, significantly improving source localization accuracy for LR EEG. Overall, the method not only enhances EEG spatial resolution but also supports high-precision EEG imaging analysis in clinical and research applications.
@@ 4 Discussion and Conclusion
Experimental results show that while the proposed STADMs outperform existing methods in reconstruction performance, larger scaling factors reveal a significant difference between the synthetic SR EEG and the ground truth, as illustrated in Figure 3. We identified several possible reasons for this: i) At larger scaling factors, the available spatial information for the model is limited, making it difficult to accurately capture the spatial relationships between channels; ii) The inherently low SNR of EEG increases the difficulty of reconstruction for the model; iii) Significant individual differences between subjects make it challenging for the model to uniformly learn the mapping between HR EEG and LR EEG across different subjects. To address the issue of individual differences, personalized strategies can be designed, which is one of our future research priorities. Despite some limitations, the proposed method has demonstrated its effectiveness in downstream tasks and offers a new approach for enhancing the spatial resolution of other physiological signals. In the future, we will further optimize this framework and apply it to practical clinical scenarios in epilepsy diagnosis, aiding doctors in the preoperative assessment for epilepsy focus resection surgery.